OpenStudy (anonymous):

Im confused I'm not sure what to do first. 5-c/3=2c+2/-4

OpenStudy (anonymous):

Pemdas, what do you see ehre according to this rule? which one should you add or subtract to the other side?

OpenStudy (anonymous):

you should subtract c and 2

OpenStudy (anonymous):

c/3 is division, also 2/-4.... in this equation, what you have to do is make the terms different in both sides. so yes.

OpenStudy (anonymous):

that gives us 5/3= 3c + 2 / -4

OpenStudy (anonymous):

remember, when you have a subtraction sign (left on this case) and you want to take that away, you have to ADD it to the other term in the other side, giving us 3 c... Correct me if I am wrong, if im not, what then?

OpenStudy (anonymous):

brb in 2 getting paper....do you understand this? :)

OpenStudy (anonymous):

Now what do I do first with 5/3= 3c + 2 / -4

OpenStudy (anonymous):

Hello are you still there

OpenStudy (anonymous):

ok, if we're doing this right, now we have to add the rest of the right equation, so 2, and /-4. ok, 2 MINUS 5 is 3, now, TIMES -4 is -12/3=3c

OpenStudy (anonymous):

yes! now, One thing I noticed and that troubled me was the division...to "balance" out the equation and finx X you have to add the real numbers, or X even to one side....however, when I multiplied -4 to get it to the otehr side I wasnt sure which one to multiply by....

OpenStudy (anonymous):

I mean, you might even have to multiply the whole left side, like (-4) 5/3= 3c + 2 / -4

OpenStudy (anonymous):

what do you think? :(

OpenStudy (anonymous):

now that I tihnk about it, it might give you the answer if you follow pemdas correctly, any how, lets finish this.

OpenStudy (anonymous):

-12/3=3c you have to solve this last part now, what is -12 divided by 3?

OpenStudy (anonymous):

Ok I'm still confused about this part 2, and /-4. ok, 2 MINUS 5 is 3, now, TIMES -4 is -12/3=3c

OpenStudy (anonymous):

hello are you still there

OpenStudy (anonymous):

ok, so, in the right side we had to add...= 3c + 2 / -4....see the two? to add it to the other side, we have to subtract it to the real number on the left, and we dont subtract the number that the left is being divided by either {5/3}

OpenStudy (anonymous):

yes, you can know that im still here because it says im typing a reply.

OpenStudy (anonymous):

with the 2 what real number do we have to subtract it to

OpenStudy (anonymous):

now, we have to subtract because they are both constituents of each other, meaning they make the same thing, however, for them to do that you have to solve it by knowing that if you did this to one side....so you subtracted 2...however, for the otehr side to be complete, as in reverse the process of the actual equation to make it true, and then find the variable by division or multiplication

OpenStudy (anonymous):

erm, above I TRIED explaining you what this whole thing is....any how, you subtract 2 by 5, because you cant subtract 2 by what you are dividing.... however, you can multiply what you are dividing....ahh....

OpenStudy (anonymous):

so, 5/3= 3c + 2 / -4, the plus two, you subtract it to the other real number variable that is not division or multiplication, or exponents...

OpenStudy (anonymous):

so, you get 3....

OpenStudy (anonymous):

Ok so 2-5 is -3 right?

OpenStudy (anonymous):

now, for /4, you multiply by the number that is being divided, seems confusing but remmember, the two sides are mirrors of each other, one thing has a relationshop with the other thing...

OpenStudy (anonymous):

wow....um....that confused me alot when I was in your grade....hmm,....yes, It is that because I did it the wrong way possibly, however, it is still correct and our answer is correct...

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

ah! ok, so, msot of the time the questions you have to follow quite obiously in a simple manner...

OpenStudy (anonymous):

look, if you had ( and this is hard to explain because in these types of formulas, THEY HAVE TO BE TRUE!!1) 2x-3=5x+3, would you subtract the 3 or add the three on the other side?

OpenStudy (anonymous):

(oh, and by the way, this question is untrue because 2x DOES NOT = 3X.

OpenStudy (anonymous):

oh, and it doesnt have to be true, like I just showed.

OpenStudy (anonymous):

another example is 2x+3=2x+4, if you choose to subtract the 4 to make the other side equal, you would get -1. and this is still correct, you can solve it this way.

OpenStudy (anonymous):

want to do one of these problems together? :)

OpenStudy (anonymous):

I'm not trying to be mean but I'm still confused about how to solve this problem my teacher really never taught us how to do these type of problems so that's why I'm on open study. com looking for help and the way your teaching me is making me more confused. sorry :( :( :(

OpenStudy (anonymous):

Ok, it's ok, this will always be somewhat confusing, but you were correct, it's nothing really it's just how you solve it.

OpenStudy (anonymous):

so, 2x-3=6x-1.

OpenStudy (anonymous):

thanks for understanding i used to like math but ever since i've gotten deeper into algebra 1 it's become more difficult

OpenStudy (anonymous):

solve either side by removing one of the variables, lets ADD 3 to 1. giving us 2, because 3-1 is 2. Whenever you ADD something to has a MINUS sign you subtract that & that.

OpenStudy (anonymous):

ok, so that gives us 2x=6x-2

OpenStudy (anonymous):

now, we can STILL add 2x to 6x, or add 6x to 2x, its really the same thing, your going to have to end up dividing any way

OpenStudy (anonymous):

so we get 8x=2

OpenStudy (anonymous):

what is 8x/2 is .25, x= .25 MAN I really hope your not confused XD

OpenStudy (anonymous):

what im trying to show you is it doesn't matter which side you remove or add the variable by.

OpenStudy (anonymous):

oh, and a variable is actually X, the other number's that are part of the equation to FIND the variable are REAL numbers.

OpenStudy (anonymous):

and also, please dont try to hard, if you dont understand it, and dont take a break, you will fall into a crux were you can't really pay attention or learn anything....do alot of things to avoid this, during class look around the room, in here, EAT, eating help me ALOT to to pay attention, and keep an attentive posture, that is important to.

OpenStudy (anonymous):

ill be answering others question, just reply me and I'll be back :)